0.95 2. 100 points] for 60 students in hy fall 2013 STA 215 sections is: У- 38.831 + 1.399x . Find the residual for a studentCan someone explain how to do it
Answer #1

Given regression equation: y = 38.831 + 1.399x

At (x, y) = (80, 140):

Predicted y = 38.831 + 1.399*80

Predicted y = 150.75

Hence,

Residual = Actual - Predicted = 140 - 150.75 = -10.75

Option D is correct.